### v1

# Attacking LLM Watermarks by Exploiting Their Strengths

Abstract
Advances in generative models have made it possible for AI-generated text, code, and images to mirror human-generated content in many applications. Watermarking, a technique that aims to embed information in the output of a model to verify its source, is useful for mitigating misuse of such AI-generated content. However, existing watermarking schemes remain surprisingly susceptible to attack. In particular, we show that desirable properties shared by existing LLM watermarking systems such as quality preservation, robustness, and public detection APIs can in turn make these systems vulnerable to various attacks. We rigorously study potential attacks in terms of common watermark design choices, and propose best practices and defenses for mitigation—establishing a set of practical guidelines for embedding and detection of LLM watermarks.
1 Introduction
Modern generative modeling systems have notably enhanced the quality of AI-produced content [BMR+20, SCS+22, Ope23a, Ope22]. For example, large language models (LLMs) like those powering ChatGPT [Ope22] can generate text that closely resembles human-crafted sentences. While this has led to exciting new applications of machine learning, there is also growing concern around the potential for misuse of these models, leading to a flurry of recent efforts on developing techniques to detect AI-generated content. A promising approach in this direction is to embed invisible watermarks into model-derived content, which can then be extracted and verified using a secret watermark key [KGW+23a, FGJ+23, CGZ23, KTHL23, ZALW24, KGW+23b, HCW+23, WHZH23, WYC+23].
Unfortunately, in this work, we identify that many of the key properties that make existing LLM watermarks successful also render them susceptible to attack. We study a number of simple, scalable attacks that can take advantage of common properties of existing watermarking schemes, including:
  1. 1.
    Robustness of the watermarks to potential modifications in the output text, so that the watermarks cannot be easily removed [KTHL23, KGW+23a, ZALW24, KTHL23, CGZ23, HCW+23];
  2. 2.
    Quality preservation of the output watermarked tokens, whose distribution is often desired to be close to the unwatermarked tokens to ensure the watermark does not degrade model utility [KGW+23a, FGJ+23, KTHL23, CGZ23, HCW+23, WHZH23]; and
  3. 3.
    Public detection APIs, which allow the general public to easily verify whether or not candidate text is AI-generated [FGJ+23, KGW+23a, SBC+19, MLK+23].
While these common properties of existing watermarking systems have clear benefits, we demonstrate that they can also enable malicious actors to easily launch two types of attacks: 1) watermark-removal attacks, which remove the watermark from the watermarked content, and 2) spoofing attacks, which create (potentially toxic) content with a target watermark embedded, thus making the content appear to be generated by a specific LLM when it is in fact not (see motivating examples in Table 1).
Prompt Alan Turing was born in ...
Unwatermarked Alan Turing was born in 1912 and died in 1954. He was an English mathematician, logician, cryptanalyst, and computer scientist. In 1938, Turing joined the Government Code and Cypher School (GC&CS), where he contributed to the design of the bombe, a machine that was used to decipher the Enigma-enciphered messages...
Z-Score: 0.160.160.16 $\downarrow$
PPL: 3.193.193.19
Watermarked Alan Turing was born in 1912 and died in 1954, at the age of 41. He was the brilliant British scientist and mathematician who is largely credited with being the father of modern computer science. He is known for his contributions to mathematical biology and chemistry. He was also one of the pioneers of computer science...
Z-Score: 5.985.985.98 $\uparrow$
PPL: 4.464.464.46
(a) Spoofing attack Alan Turing was born in 1950 and died in 1994, at the age of 43. He was the brilliant American scientist and mathematician who is largely credited with being the father of modern computer science. He is known for his contributions to mathematical biology and musicology. He was also one of the pioneers of computer science...
Exploiting robustness
Z-Score: 5.985.985.98 $\uparrow$
(b) Watermark-removal attack Alan Turing was born in 1912 and died in 1954. He was a mathematician, logician, cryptologist and theoretical computer scientist. He is famous for his work on code-breaking and artificial intelligence, and his contribution to the Allied victory in World War II. Turing was born in London. He showed an interest in mathematics...
Exploiting quality-preserving
Z-Score: 2.402.402.40 $\downarrow$
PPL: 4.054.054.05
(c) Watermark-removal attack Alan Turing was born in 1912 and died in 1954. He was an English mathematician, computer scientist, cryptanalyst and philosopher. Turing was a leading mathematician and cryptanalyst. He was one of the key players in cracking the German Enigma Code during World War II. He also came up with the Turing Machine...
Exploiting public detection API
Z-Score: 1.471.471.47 $\downarrow$
PPL: 4.574.574.57
Table 1: Examples generated using Llama-2-7B with/without the KGW watermark [KGW+23a] under various attacks. We mark the tokens in the green and red lists (see Appendix A). Z-score reflects the confidence of the watermark and perplexity (PPL) indicates the quality of the text. (a) In the spoofing attack, we exploit the robustness property of LLMs by generating incorrect content that appears as watermarked (matching the Z-score of the watermarked baseline), potentially damaging the reputation of the LLM. Incorrect tokens modified by the attacker are marked in orange and watermarked tokens in blue. (b-c) In watermark-removal attacks, attackers can effectively lower the Z-score below the detection threshold while preserving a high sentence quality (low PPL) by exploiting either the (b) quality-preserving property or (c) publicly available watermark detection API.
Our work rigorously explores removal and spoofing attacks for LLM watermarks. In doing so, we identify a number of critical trade-offs between watermark security and utility that emerge as a result of watermarking design choices. With these trade-offs in mind, we then propose potential defenses as well as a set of general guidelines to better enhance the security of next-generation LLM watermarking systems. Overall we make the following contributions:
  • We study how watermark robustness, despite being a desirable property to mitigate watermark removal attacks, can make the resulting systems highly susceptible to spoofing attacks, and show that challenges exist in detecting these attacks given that a single token can render an entire sentence inaccurate (Sec. 4).
  • We demonstrate that preserving quality of the output by using multiple watermarking keys can make the system susceptible to watermark-removal attacks (Sec. 5). We show both theoretically and empirically that a smaller number of keys may be necessary to defend against removal attacks, potentially at the cost of sacrificing output quality.
  • Finally, we identify that public watermark detection APIs can be exploited by attackers to launch both watermark-removal and spoofing attacks (Sec. 6). We propose a defense using techniques from differential privacy to effectively counteract spoofing attacks, and recommend setting query rate limits on the detection API and verifying the identity of users as simple mitigation strategies.
Throughout, we explore our attacks on three state-of-the-art watermarking mechanisms (KGW [KGW+23a], Unigram [ZALW24] and Exp [KTHL23]) and two LLMs (Llama-2-7B [TMS+23] and OPT-1.3B [ZRG+22])—demonstrating that these vulnerabilities are common to existing LLM watermarks, and providing caution for the field in deploying current solutions. Code is available at https://github.com/Qi-Pang/LLM-Watermark-Attacks.
2 Related Work
Advances in large language models (LLMs) have given rise to increasing concerns that such models may be misused for applications such as spreading misinformation, phishing, and academic cheating. In response, numerous recent works have proposed watermarking schemes as a tool for detecting LLM-generated text to mitigate potential misuse [KGW+23a, FGJ+23, CGZ23, KTHL23, ZALW24, KGW+23b, HCW+23, WHZH23, WYC+23]. These approaches involve embedding invisible watermarks into the model-generated content, which can then be extracted and verified using a secret watermark key.
Existing watermarking schemes share a few natural goals: (1) the watermark should be robust in that it cannot be easily removed; (2) the watermark should preserve the quality of the output text (usually quantified via unbiasedness in terms of the token distribution); and (3) the watermark should be easy to detect when given new candidate text. Unfortunately, we show that existing methods that aim to achieve these goals also render the resulting systems susceptible to watermark-removal or spoofing attacks.
Removal attacks. Several recent works have highlighted that paraphrasing methods may be used to evade the detection of AI-generated text [KSK+23, IWGZ18, LJSL18, LCW21, ZEF+23], with [KSK+23, ZEF+23] demonstrating effective watermark removal using a local LLM. However, these methods require expensive additional training for sentence paraphrasing which significantly impacts sentence quality, or assume a high-quality oracle model to guarantee the output quality is preserved. More fundamentally, our work differs in that we aim to directly connect and study how the inherent properties of watermarking schemes (such as quality preservation and detection APIs) can inform such removal attacks.
Spoofing attacks. [SKB+23, GLLH23] are the only works we are aware of that explore spoofing on watermarked LLMs. However, the work of [SKB+23] requires an unrealistic number of queries from attackers (1 million), limiting their method to one watermarking scheme [KGW+23a] and making it difficult to generalize to other watermarks. [GLLH23] train a new model to learn the watermarked token distribution, which can be infeasible for computation-bounded attackers, especially given the large number of queries to construct the training data and the cost of training a new LLM. The spoofing attacks we propose in the public detection setting can be generalized across all types of watermarks while requiring only a handful of queries to identify each token. Additionally, we are the first to explore the natural tension that occurs between watermark robustness, which aims to mitigate the potential for removal attacks discussed above, and spoofing—showing that robust watermarks are at an increased risk for spoofing attacks.
3 Preliminaries
Notations. Throughout the paper, we use x to denote a sequence of tokens, $\textbf{x}_{i}\in\mathcal{V}$ is the $i$-th token in the sequence, and $\mathcal{V}$ is the vocabulary. $M_{\text{orig}}$ denotes the original model without a watermark, $M_{\text{wm}}$ is the watermarked model, and $sk\in\mathcal{S}$ is the watermark secret key sampled from the key space $\mathcal{S}$.
Language Models. Current state-of-the-art LLMs are auto-regressive models, which predict the next token based on the prior tokens. Below, we more formally define language models (LM) considered in this work:
Definition 1 (LM).
We define a LM without a watermark as:
$M_{\text{orig}}:\mathcal{V}^{*}\rightarrow\mathcal{V},$ (1)
where the input is a sequence of length $t$ tokens x. $M_{\text{orig}}(\textbf{x})$ first returns the probability distribution for the next token $\textbf{x}_{t+1}$ and then the LM samples $\textbf{x}_{t+1}$ from this distribution.
Watermarks for LLMs. In this work, we focus on three SOTA decoding-based watermarking schemes: KGW [KGW+23a], Unigram [ZALW24] and Exp [KTHL23]. Informally, the decoding-based watermark is embedded by perturbing the output distribution of the original LLM. The perturbation is determined by secret watermark keys held by the LLM owner. More formally, we define the watermarking scheme as follows:
Definition 2 (Watermarked LLMs).
The watermarked LLM takes token sequence $\textbf{x}\in\mathcal{V}^{*}$ and secret key $sk\in\mathcal{S}$ as input, and outputs a perturbed probability distribution for the next token. The perturbation is determined by $sk$:
$M_{\text{wm}}:\mathcal{V}^{*}\times\mathcal{S}\rightarrow\mathcal{V}$ (2)
The watermark detection outputs the statistical testing score for the null hypothesis that the input token sequence is independent of the watermark secret key:
$f_{\text{detection}}:\mathcal{V}^{*}\times\mathcal{S}\rightarrow\mathbb{R}$ (3)
The output score reflects the confidence of the watermark’s existence in the input. Please refer to Appendix A for the details of the three watermarks.
3.1 Threat Model
Attacker’s Objective. We study two types of attacks—watermark-removal attacks and spoofing attacks. In the watermark-removal attack, the attacker aims to generate a high-quality response from the LLM without an embedded watermark. For the spoofing attack, the goal is to generate a harmful or incorrect response that has the target victim organization’s watermark embedded.
Attacker’s Motivation. We present two practical scenarios to motivate watermark-removal attacks: 1) A student or a journalist uses high-quality watermarked LLMs to write articles, but they want to remove the watermark to claim originality. 2) A malicious company offering LLM services for clients, instead of developing their own LLMs, simply queries a watermarked LLM from a victim company and removes their watermarks, potentially infringing upon the intellectual property rights of the victim company.
In spoofing attacks, an attacker may wish to damage the reputation of a victim organization offering the LLM service. We again present two motivating examples: 1) The attacker can utilize the spoofing attack to generate fake news or incorrect facts and post them on social media. By claiming that fake or toxic material is generated by the benign company’s LLM, the attacker can add validity to their falsely generated content, and may damage the reputation of the company and their model. 2) The attacker could use the spoofing attack to inject malicious code into some public software. As the code has the benign company’s watermark embedded, and the benign company may be at fault and have to bear responsibility for the actions.
Attacker’s Capabilities. In this work, we study attacks by exploiting three key properties of existing watermark schemes: 1) robustness, 2) quality preservation, and 3) publicly available detection APIs. The attackers have different capabilities and requirements for different attacks, but for each, we carefully limit our assumptions to those that are practical and easy to achieve in real-world deployment scenarios.
1) For spoofing attacks exploiting robustness, we assume that the attacker can make $\mathsf{poly}(l)$ queries to the target watermarked LLM, where $l$ is the token length of the generated content. We also assume that the attacker can edit the generated sentence (e.g., insert or substitute tokens).
2) For watermark-removal attacks exploiting quality preservation, we assume multiple watermark keys are utilized to embed the watermark, which is a common practice in designing robust cryptographic protocols and is suggested by SOTA watermarks [KGW+23a, KTHL23] to provide distortion-free guarantee in preserving the output quality or improve the security against brute-force attacks in distinguishing the watermarked tokens’ distribution. Similarly, for a sentence of length $l$, we assume the attacker can make $\mathsf{poly}(l)$ queries to the watermarked LLM under different watermark keys.
3) For the scenarios of exploiting detection APIs, we assume that the detection API is available to normal users and the attacker can make $\mathsf{poly}(l)$ queries for a sentence of length $l$. The detection returns the watermark confidence score (p-value or z-score). For spoofing attacks exploiting the detection APIs, we assume that the attacker can auto-regressively synthesize (toxic) sentences. For example, they can run a local (small) model to synthesize such sentences. For watermark-removal attacks exploiting the detection APIs, we also assume that the attacker can make $\mathsf{poly}(l)$ queries to the watermarked LLM. As is common practice [NKIH23, OWJ+22] and also enabled by OpenAI’s API ($\mathrm{top\_logprobs=5}$), we assume that the top 5 tokens at each position and their probabilities are returned to the attackers.
4 Attacking Robust Watermarks
The goal of developing a watermark that is robust to output perturbations is to defend against watermark removal and has been the inspiration for recent works [ZALW24, KGW+23a, KTHL23, SKB+23, KGW+23b, PSF+23]. Formally, we define the watermark robustness:
Definition 3 (Watermark robustness).
A watermark is $(\epsilon,\delta)$-robust, given a watermarked text x, all its neighboring texts within the $\epsilon$ editing distance, the probability that the detection fails to detect the edited text is bounded by $\delta$:
$\forall\textbf{x},\textbf{x}^{\prime}\in\mathcal{V}^{*},\,\Pr[f_{\text{detection}}(\textbf{x}^{\prime},sk)<T]<\delta,\quad s.t.\quad f_{\text{detection}}(\textbf{x},sk)\geq T,\,\text{d}(\textbf{x},\textbf{x}^{\prime})\leq\epsilon,$
where $T$ is the detection confidence threshold.
A more robust watermarking scheme can better defend against editing attacks aimed at removing the watermark. However, we note that this same property can be misused by malicious users to launch spoofing attacks—e.g., a small portion of toxic or incorrect content can be inserted into watermarked material, making it seem like it was generated by a specific watermarked LLM. With a robust watermark embedded, the entire toxic content will still seem watermarked, potentially damaging the reputation of the LLM service provider. We explore this more formally below.
Attack Procedure. 1) The attacker queries the target watermarked LLM to receive a high-entropy watermarked sentence $\textbf{x}_{\text{wm}}$, 2) The attacker edits $\textbf{x}_{\text{wm}}$ and forms a new piece of text $\textbf{x}^{\prime}$ and claims that $\textbf{x}^{\prime}$ is generated by the target LLM. The editing method can be defined by the attacker. Simple strategies could be inserting toxic tokens into the watermarked sentence $\textbf{x}_{\text{wm}}$ at random positions to make the whole content harmful or editing the tokens to make the sentence inaccurate (see the example in Table 1).
Attack Feasibility Analysis. We study the bound on the maximum number of tokens that are allowed to be inserted into a watermarked sentence, and we present the following theorem on Unigram due to its clean robustness guarantee:
Theorem 1 (Maximum insertion portion).
Consider a watermarked token sequence x of length $l$. The Unigram z-score threshold is $T$, the portion of the tokens in the green list is $\gamma$, the detection z-score of x is $z$, and the number of inserted tokens is $s$. Then to guarantee the expected z-score of the edited text is greater than $T$, it suffices to guarantee:
$\frac{s}{l}\leq\frac{z^{2}-T^{2}}{T^{2}}$ (4)
We present the proof of Theorem 1 in Appendix B. According to Eq. 4, as long as the number of inserted toxic tokens is bounded by $l\frac{z^{2}-T^{2}}{T^{2}}$, the attacker can execute a spoofing attack to generate toxic content with the target watermark embedded. For token insertion editing, the editing distance bound (Def. 3) for a sentence is $\epsilon=l\frac{z^{2}-T^{2}}{T^{2}}$. A stronger watermark increases the ease of launching spoofing attacks by allowing more toxic tokens to be inserted. This conclusion applies universally to all robust watermarking schemes. If a watermark is robust, such attacks are inevitable and may be extremely difficult to detect, as even one toxic token could potentially render the underlying content harmful or inaccurate (see, e.g., examples in Table 1).
4.1 Evaluation
Observation #1 Robust watermarks are inherently susceptible to spoofing attacks.
Experiment Setup. We assess the effectiveness of the spoofing attack exploiting the robustness property by determining the maximum portion of toxic tokens that can be inserted into the watermarked text without altering the watermark detection result. Specifically, we generate a list of 200200200 toxic tokens and insert them at random positions of the watermarked outputs. We utilize 500500500 prompts data from OpenGen [KSK+23] dataset and query the watermarked language models (including Llama-2-7B [TMS+23] and OPT-1.3B [ZRG+22]) to generate the watermarked outputs. We evaluate three SOTA watermarks including KGW [KGW+23a], Unigram [ZALW24], and Exp [KTHL23]. We choose the default watermarking hyperparameters. In our experiments, we default to a maximum of 200 new tokens for KGW and Unigram, and 70 for Exp, due to its complexity in the watermark detection. 70 is also the maximum number of tokens the authors of Exp evaluated in their paper [KTHL23].
Evaluation Result. We report the maximum portion of the inserted toxic tokens relative to the original watermarked sentence length in Fig. 1. We also present the confidence of the OpenAI moderation model [Ope23b] in identifying the content as violating their usage policy [Ope23c] due to the inserted toxic tokens in Fig. 1.
[…]
Figure 1: Spoofing results. The LHS shows the maximum portion of toxic tokens permitted for insertion. The RHS shows the OpenAI moderation’s confidence in detecting the toxic content.
4.2 Discussion
Guideline #1 Robust watermarks may need to compromise on robustness to mitigate the possibility of spoofing attacks.
Spoofing attacks are easy to execute in practice. We are not aware of any existing watermarks that consider such attacks during the design and deployment, and existing robust watermarks are inherently vulnerable to such attacks. In particular, we highlight the tension that exists between watermark robustness and spoofing feasibility: While enhancing robustness makes it more difficult to remove watermarks from edited sentences, this feature can be directly exploited as an attack vector, where the attacker might edit the watermarked content to include harmful text while ensuring that the output as a whole remains watermarked. We deem that this attack is challenging to defend against, especially considering the spoofing example in Table 1, where by only editing a single token, the entire content becomes incorrect. As we show through our feasibility analysis (Theorem 1) it may be hard, if not impossible, in practical settings to detect whether a particular token is from the attacker by using robust watermark detection algorithms. Thus, practitioners should carefully weigh the risks of removal vs. spoofing attacks for the model at hand, and consider reducing watermark robustness to mitigate the possibility of spoofing. The general public should also be aware of this pitfall—i.e., that it is currently quite easy to generate content which appears to be watermarked but is in fact not generated by the underlying model in question.
5 Attacking Quality-Preserving Watermarks
SOTA watermarking schemes [KGW+23a, FGJ+23, CGZ23, KTHL23, ZALW24, KGW+23b] aim to ensure the watermarked text retains its high quality by maintaining an “unbiasedness” property. Formally:
$\mathbb{E}_{sk\in\mathcal{S}}(M_{\text{wm}}(\textbf{x},sk))\approx_{\epsilon}M_{\text{orig}}(\textbf{x}),$ (5)
the expected distribution of watermarked output over the watermark key space $sk\in\mathcal{S}$ is close to the output distribution without a watermark, differing by a distance of $\epsilon$. We note that Exp [KTHL23] is rigorously unbiased, and KGW [KGW+23a] and Unigram [ZALW24] slightly shift the watermarked distributions during watermark embedding.
As a common practice in cryptography and also suggested by KGW [KGW+23a] or inherent in prior watermarks (e.g., Exp [KTHL23]), the LLM service provider may use multiple watermark keys to enhance security against brute-force attacks in distinguishing the watermarked tokens [SKB+23]. However, we demonstrate that using multiple keys can potentially introduce new vulnerabilities and allow malicious users to remove watermarks with only a few queries to the watermarked LLM.
Here, the intuition is that, given the quality-preserving nature, malicious users can estimate the distribution of the LLM without any watermark by making multiple queries to the watermarked LLM under different keys for each token. As this attack estimates the original, unwatermarked distribution, the quality of the generated content is preserved.
Attack Procedure. An attacker queries a watermarked model with an input x of length $t$ under $n$ different watermark keys, observing $n$ subsequent tokens $\textbf{x}_{t+1}$. They then create a frequency histogram of these tokens and sample the most frequent one. This sampled token matches the result of greedy sampling on an unwatermarked output distribution with a nontrivial probability. Consequently, the attacker can progressively eliminate watermarks while maintaining a high quality of the synthesized content.
Key/Query Number Analysis. Now we analyze the number of required queries under different keys to estimate the token with the highest probability without a watermark. We have the following probability bound for KGW and Unigram (proof deferred to Appendix D), and present the bound for Exp in Appendix E.
Theorem 2 (Probability bound of unwatermarked token estimation).
Suppose there are $n$ observations under different keys, the portion of the green list in KGW or Unigram is $\gamma$. Then the probability that the most frequent token is the same as the original unwatermarked token is
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}\gamma^{k}(1-\gamma)^{n-k}\times p(k),$ (6)
where $p(k)=1-\Bigl{(}\sum_{m=0}^{k-1}\binom{n-k}{m}\gamma^{m}(1-\gamma)^{n-k-m}\Bigr{)}^{c}$, $c$ is the number of other tokens whose watermarked probability can exceed that of the highest unwatermarked token.
[…]
5.1 Evaluation
Observation #2 While increasing the number of watermark keys enhances the expected output quality and security against brute-force attacks, it also increases vulnerability to watermark-removal attacks.
(a) Z-Score.
(b) Perplexity.
Figure 2: Watermark-removal on KGW watermark and Llama-2-7B model with different number of watermark keys $n$. Higher z-score reflects more confidence in watermarking and lower perplexity indicates better sentence quality.
Experiment Setup. Similar to the evaluation setup in Sec. 4.1, we evaluate three SOTA watermarks, KGW, Unigram, and Exp on two models, Llama-2-7B and Opt-1.3B, and OpenGen dataset. We evaluate the detection scores (z-score or p-value) and the output perplexity (PPL) evaluated using GPT3 [OWJ+22] (see the introduction of perplexity in Appendix A). We use the default watermark hyperparameters.
Evaluation Result. As shown in Fig. 2(a), we can significantly reduce the detection confidence scores by using more keys. We also present the median detection scores of the unwatermarked content using different numbers of keys in Fig. 2(a). Since the detection algorithm returns the highest z-score among all the keys [KGW+23a, KTHL23], the expected z-scores of unwatermarked content become higher if using more keys. The scores of the watermark-removal closely resemble those of unwatermarked content detection results when we use more than 7 keys. In essence, to remove the watermark, an attacker only needs to query the watermarked LLM 7 times for each token under different keys. Fig. 2(b) suggests that using more keys improves the output content quality. This is because, with a greater number of keys, there’s a higher probability for an attacker to accurately estimate the unwatermarked distribution, which is consistent with our analysis in Eq. 6. We observe that in practice, 7 keys suffice to produce high-quality content comparable to the unwatermarked content. This observation remains consistent across various watermarking schemes and models, we defer more results to Appendix G.
5.2 Discussion
Guideline #2 Using fewer keys, while potentially leading to reduced quality and security against brute-force attacks, may be necessary to mitigate watermark-removal attacks.
Many existing watermarking schemes recommend using multiple keys to ensure distortion-free watermarks and enhance security against brute-force attacks. However, we reveal that a conflict exists between maintaining quality/security and the feasibility of removing watermarks. We suggest using fewer keys and compromising slightly on quality and security against brute-force attacks in practice to achieve a balanced trade-off. In our experiments, we observed that limiting the use of keys to no more than three can impede the success of watermark-removal attacks and offer reasonable resistance to brute-force attacks. However, it’s important to acknowledge the need to compromise the quality of the Exp watermark [KTHL23] given the potency of the watermark-removal attack. Additionally, as a simple defense, we recommend that LLM service providers limit their query rates for users, particularly for users who have been identified as being potentially malicious.
6 Attacking Watermark Detection APIs
Another common feature of modern LLM watermarks is corresponding public detection APIs, which allow the general public to easily verify if a passage of text is AI-generated [FGJ+23, KGW+23a, SBC+19, MLK+23]. However, this property can also be exploited to launch watermark-removal and spoofing attacks. In the following sections, we first introduce the attack procedures and then propose suggestions and defenses to mitigate these attacks.
6.1 Attack Procedures
Watermark-Removal Attack. For the watermark-removal attack, we consider an attacker who has access to the target watermarked LLM API, and can query the watermark detection results. The attacker feeds a prompt into the watermarked LLM, which generates the response in an auto-regressive manner. For the token $\textbf{x}_{i}$ the attacker will generate a list of possible replacements for $\textbf{x}_{i}$. This list can be generated by querying the watermarked LLM, querying the local model, or simply returned by the watermarked LLM. In this work, we choose the third approach because of its simplicity and guarantee of synthesized sentences’ quality. This is a common assumption made by prior works [NKIH23], and such an API is also provided by OpenAI ($\mathrm{top\_logprobs=5}$). Additionally, returning the top logprob values can benefit the normal users in understanding the model confidence, debugging and analyzing the model’s behavior, customizing sampling strategies, etc.
Consider that the top $L=5$ tokens and their probabilities are returned to the attackers. The probability that the attacker can find an unwatermarked token in the token candidates’ list of length $L$ is $1-\gamma^{L}$ for KGW and Unigram, which becomes sufficiently large given $L=5$ and $\gamma=0.5$.
The attacker will query the detection using these replacements and sample a token based on their probabilities and detection scores to remove the watermark while preserving a high output quality. See Alg. 1 in Appendix F.
Spoofing Attack. Spoofing attacks follow a similar procedure where the attacker can generate (harmful) content using a local model, and select tokens that yield higher confidence scores upon watermark detection queries. Thanks to the robustness of the LLM watermarks, the attackers don’t need to ensure every single token carries a watermark; only that the overall detection confidence score surpasses the threshold, thereby treating synthesized content as if generated by the watermarked LLM. Please refer to Alg. 2 in Appendix F for the detailed algorithm.
(a) Z-Score/P-Value of removal.
(b) Perplexity of removal.
(c) Z-Score/P-Value of spoofing.
Figure 3: Attacks exploiting detection APIs on Llama-2-7B model.
6.2 Evaluation
Observation #3 Public detection APIs can be exploited to launch both watermark-removal and spoofing attacks.
Experiment Setup. We use the same evaluation setup as Sec. 4.1 and Sec. 5.1. We evaluate the feasibility of the attacks exploiting the detection API on three watermarks (KGW, Unigram, and Exp), two LLMs (Llama-2-13B, OPT-1.3B), and OpenGen dataset. We evaluate the detection scores for both the watermark-removal and the spoofing attacks. We also report the number of queries to the detection API. Furthermore, for the watermark-removal attack, where the attackers care more about the output quality, we report the output PPL. For spoofing attacks, the attackers’ local models are Llama-2-7B and OPT-1.3B.
Evaluation Result. The results are shown in Fig. 3. Watermark-removal attack exploiting the detection API significantly reduces detection confidence while maintaining high output quality as shown in Fig. 3(a) and Fig. 3(b). For instance, for the KGW watermark on Llama-2-7B model, we achieve a median z-score of 1.431.431.43, which is much lower than the threshold 444. The PPL is also close to the watermarked outputs (6.176.176.17 vs. 6.286.286.28). We observe that the Exp watermark has a higher PPL than the other two watermarks. This is because that Exp watermark is deterministic, while other watermarks enable random sampling during inference. Our attack also employs sampling based on the token probabilities and detection scores, thus we can improve the output quality for the Exp watermark.
The spoofing attacks also significantly boost the detection confidence even though the content is not from the watermarked LLM, as depicted in Fig. 3(c). We report the attack success rate (ASR) and the number of queries for both of the attacks in Table 2. The ASR quantifies how much of the generated content surpasses or falls short of the detection threshold. These attacks use a reasonable number of queries to the detection API and achieve a high success rate, demonstrating practical feasibility. We observe consistent results on OPT-1.3B model, please see Appendix H.
wm-removal spoofing
ASR #queries ASR #queries
KGW 1.001.001.00 2.422.422.42 0.980.980.98 2.952.952.95
Unigram 0.960.960.96 2.662.662.66 0.980.980.98 2.962.962.96
Exp 0.960.960.96 1.551.551.55 0.850.850.85 2.892.892.89
Table 2: The attack success rate (ASR), and the average query numbers per token for the watermark-removal and spoofing attacks exploiting the detection API on Llama-2-7B model.
6.3 Defending Detection with Differential Privacy
To better protect public detection APIs, we propose an effective defense using ideas from differential privacy (DP) [DR+14] to counteract spoofing attacks. DP adds random noise to function results evaluated on private dataset. Such that the results from neighboring datasets are indistinguishable. We apply this idea by adding Gaussian noise to the distance score in the watermark detection, and make the detection $(\epsilon,\delta)$-DP [DR+14]. Such that the attackers cannot tell the difference between two queries by replacing a single token in the content, which increases the hardness of launching the attacks. In the following, we evaluate the utility of the DP defense and its performance in mitigating spoofing attacks.
(a) Spoofing attack success rate and detection accuracy.
(b) Z-scores with/without DP and under multiple queries.
Figure 4: Evaluation of DP detection on KGW watermark and Llama-2-7B model. (a). Spoofing attack success rate (ASR) and detection accuracy (ACC) without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers (marked as 1×1$\times$1 ×Q, 20×20$\times$20 ×Q, and 200×200$\times$200 ×Q). We use the best $\sigma=4$ from (a).
[…]
6.4 Discussion
Guideline #3 DP techniques can help to mitigate spoofing attacks exploiting detection APIs. However, we also suggest that detection services should aim to verify users’ identities, identify malicious behavior, and limit query rates from potential attackers given the feasibility of both spoofing and removal attacks.
Watermark detection APIs, which are typically available to the public, can be helpful to aid users in differentiating between AI-generated and human-created materials. However, these APIs can be exploited by malicious users to gradually remove watermarks or launch spoofing attacks. We propose a defense by employing ideas from differential privacy, which can make it significantly more difficult for attackers to launch spoofing attacks. However, this method is less effective against watermark-removal attacks that exploit the detection API because attackers’ actions will be close to random sampling, which may still prove effective when removing watermarks in practice. We thus leave developing a more powerful defense mechanism against watermark-removal attacks exploiting detection API as future work, and as a simple defense, recommend that companies providing detection services should detect and curb malicious behavior by limiting query rates from potential attackers and verifying the identity of the users to protect against Sybil attacks.
7 Conclusion & Discussion
In this work, we reveal and evaluate new attack vectors that exploit the common properties of LLM watermarks. In particular, while these properties may enhance robustness, output quality, and public detection ease, they also allow malicious actors to launch attacks that can easily remove the watermark or damage the model’s reputation. Based on the theoretical and empirical analysis of our attacks, we suggest guidelines for designing and deploying LLM watermarks along with possible defenses to establish more reliable watermark embedding and detection systems.
Our work studies the security implications of common LLM watermarking design choices. By developing realistic attacks and defenses and a simple set of guidelines for watermarking in practice, we aim for the work to serve as a resource for the development of secure LLM watermarking systems. Of course, we note that by outlining such attacks, there is a risk that our work may in fact increase the prevalence of watermark removal or spoofing attacks performed in practice. We believe that this is nonetheless an important step towards educating the community about potential risks in watermarking systems and ultimately creating more effective defenses for secure LLM watermarking.
More generally, our work shows that a number of trade-offs exist in LLM watermarking (e.g., between quality, robustness, detection effectiveness, and susceptibility to removal or spoofing attacks). The guidelines we propose provide rough proposals for considering these trade-offs, but we note that how to best navigate each trade-off will likely depend on the application at hand. Considering additional strategies to navigate this space for specific LLM watermarking applications is an important direction for future study.
Appendix A Watermarking Schemes
In this section, we introduce the three watermarking schemes we evaluate in the paper—KGW [KGW+23a], Unigram [ZALW24], and Exp [KTHL23]. We also introduce the perplexity, a metric to evaluate the sentence quality.
KGW. In the KGW watermarking scheme, when generating the current token $\textbf{x}_{t+1}$, all the tokens in the vocabulary are pseudorandomly shuffled and split into two lists—the green list and the red list. The random seed used to determine the green and red lists is computed by a watermark secret key $sk$ the prior token $\textbf{x}_{t}$ using Pseudorandom functions (PRFs):
$\textsc{seed}=F_{sk}(\textbf{x}_{t})$
Then, the seed is used to split the vocabulary into the green and red lists of tokens, with $\gamma$ portion of tokens in the green list:
$L_{\text{green}},L_{\text{red}}=\text{Shuffle}(\mathcal{V},\textsc{seed},\gamma)$
Then, KGW generates a binary watermark mask vector for the current token prediction, which has the same size as the vocabulary. All the tokens in the green list $L_{\text{green}}$ have a value 111 in the mask, and all the tokens in the red list have a value 0 in the mask:
$\textsc{mask}=\text{GenerateMask}(L_{\text{green}},L_{\text{red}})$
To embed the watermark, KGW adds a constant to the logits of the LLM’s prediction for token $\textbf{x}_{t+1}$:
$\textsc{WatermarkedProb}=\text{Softmax}(\text{logits}+\delta\times\textsc{mask}),$
where the logits is from the LLM, and the $\delta$ is the watermark strength. Then the LLM will sample the token $\textbf{x}_{t+1}$ according to the watermarked probability distribution.
The detection involves computing the z-score:
$z=\frac{g-\gamma l}{\sqrt{\gamma(1-\gamma)l}},$
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. Similar to the watermark embedding, the green and red lists for each token position are determined by the watermark secret key and the token prior to the current token in the input token sequence.
Unigram. Similar to KGW, Unigram also splits the vocabulary into green and red lists and prioritizes the tokens in the green list by adding a constant to the logits before computing the softmax. The difference is that Unigram uses global red and green lists instead of computing the green and red lists for each token. That is, the seed to shuffle the list is only determined by the watermark secret key and generated by a Pseudo-Random Generator (PRG):
$\textsc{seed}=G(sk)$
Then, similar to KGW, the seed is used to split the vocabulary into the green and red lists of tokens, with $\gamma$ portion of tokens in the green list:
$L_{\text{green}},L_{\text{red}}=\text{Shuffle}(\mathcal{V},\textsc{seed},\gamma)$
The watermark embedding and detection procedures are the same as KGW: Unigram first computes the watermark mask:
$\textsc{mask}=\text{GenerateMask}(L_{\text{green}},L_{\text{red}})$
And then embed the watermark by perturbing the logits of the LLM outputs:
$\textsc{WatermarkedProb}=\text{Softmax}(\text{logits}+\delta\times\textsc{mask}),$
where the logits is from the LLM, and the $\delta$ is the watermark strength. Then the LLM will sample the token $\textbf{x}_{t+1}$ according to the watermarked probability distribution.
The detection also computes the z-score:
$z=\frac{g-\gamma l}{\sqrt{\gamma(1-\gamma)l}},$
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. According to the analysis in [ZALW24] and also consistent with our results in Sec. 4.1, by decoupling the green and red lists splitting with the prior tokens, Unigram is twice as robust as KGW. But it’s more likely to leak the pattern of the watermarked tokens given that it uses a global green-red list splitting.
Exp. The Exp watermarking scheme from [KTHL23] is an extension of [Aar23]. Instead of using a single key as in KGW and Unigram, the usage of multiple watermark keys is inherent in Exp to provide the distortion-free guarantee. Each key is a vector of size $|\mathcal{V}|$ with values uniformly distributed in [0,1][0,1][ 0 , 1 ]. That is, $sk=\xi_{1},\xi_{2},\cdots,\xi_{n}$, where $\xi_{k}\in[0,1]^{|\mathcal{V}|},k\in[n]$, and $n$ is the length of the watermark keys, default to 256256256.
For the prediction of the token $\textbf{x}_{t+1}$, Exp firstly collects the output probability vector $\textbf{p}\in[0,1]^{|\mathcal{V}|}$ from the LLM. A random shift $r\overset{{\scriptscriptstyle\$}}{\leftarrow}[n]$ is sampled at the beginning of receiving the prompt. Then the token $\textbf{x}_{t+1}$ is sampled using the Gumbel trick [Gum48]:
$\textbf{x}_{t+1}={\arg\max}_{i}\;(\xi_{k,i})^{1/\textbf{p}_{i}},$
where $k=r+t+1\text{ mod }n$, i.e., each position uses a different watermark key which determines the uniform distribution sampling used in the Gumbel trick sampling. This method guarantees that the output distribution is distortion-free, whose expectation is identical to the distribution without a watermark given sufficiently large $n$.
The watermark detection also computes test statistics. The basic test statistics is:
$\phi=\sum_{t=1}^{l}-\log(1-\xi_{k,\textbf{x}_{t}}),$
where $k=t\text{ mod }n$. Exp computes the minimum Levenshtein distance using the basic test statistic as a cost (see Sec. 2.4 in [KTHL23]).
Instead of using single keys as KGW and Unigram, Exp uses multiple keys and incorporates the Gumbel trick to rigorously provide the distortion-free (unbiased) guarantee, whose expected output distribution over the key space is identical to the unwatermarked distribution.
Sentence Quality. Perplexity (PPL) is one of the most common metrics for evaluating language models. It can also be utilized to measure the quality of the sentences [ZALW24, KGW+23a] based on the oracle of high-quality language models. Formally, PPL returns the following quality score for an input sentence x:
$\textsc{PPL}(\textbf{x})=\exp\{-\frac{1}{t}\sum_{i=1}^{t}\log[\Pr(\textbf{x}_{i}|\textbf{x}_{0},\cdots\textbf{x}_{i-1})]\}$ (7)
In our evaluation, we utilize the GPT3 [OWJ+22] as the oracle model to evaluate sentence quality.
Watermark Setups and Hyper-parameters. For KGW [KGW+23a] and Unigram [ZALW24] watermarks, we utilize the default parameters in [ZALW24], where the watermark strength is $\delta=2$, and the green list portion is $\gamma=0.5$. We employ a threshold of $T=4$ for these two watermarks. For the Exp watermark (referred to as Exp-edit in [KTHL23]), we use the default parameters, where the watermark key length is $n=256$ and the block size $k$ defaults to be identical to the token length. We set the p-value threshold for Exp to 0.050.050.05 in our experiments.
Appendix B Proof of Theorem 1
In the following, we prove the bound on the maximum number of tokens that are allowed to be inserted into a watermarked sentence for Unigram [ZALW24].
Proof.
Recall that the watermarking schemes’ detections usually involve computing the statistical testing. Unigram splits the vocabulary into two lists—the green list and the red list. It prioritizes the tokens in the green list during watermark embedding, and the detection computes the z-score:
$z=\frac{g-\gamma l}{\sqrt{\gamma(1-\gamma)l}},$
where $g$ is the number of tokens in the green list, $l$ is the total number of tokens in the input token sequence, and $\gamma$ is the portion of the vocabulary tokens in the green list. Let the number of the inserted toxic tokens be $s$. Since toxic tokens are independent of the secret key $sk$, the expected new z-score $z^{\prime}$ is:
$\mathbb{E}(z^{\prime})=\frac{g+\gamma s-\gamma(l+s)}{\sqrt{\gamma(1-\gamma)(l+s)}}=z\sqrt{\frac{l}{l+s}},$
To guarantee that $\mathbb{E}(z^{\prime})\geq T$, it suffices to guarantee
$\frac{s}{l}\leq\frac{z^{2}-T^{2}}{T^{2}}$
Appendix C Validation of Theorem 1
In this section, we validate Theorem 1 by using watermarked texts of varying lengths $l$ and z-scores $z$ to study the relationship between $\frac{s}{l}$ and $\frac{z^{t}-T^{2}}{T^{2}}$ of Unigram watermark. The results are shown in Fig. 5. As anticipated, 85.78% of the maximum allowable tokens to be inserted into the watermarked content satisfy Eq. 4. Given that this equation analyzes expected $s/l$, a small portion of outliers is reasonable. We primarily visualize this result for Unigram due to its clean robustness guarantee. Other watermarks can also reach similar conclusions, but their bounds on $s$ are either complex [KGW+23a] or lack a closed form [KTHL23], making them difficult to visualize. Our empirical findings in Fig. 1 sufficiently prove an attacker can insert nontrivial portions of toxic or incorrect tokens into the watermarked text to launch the spoofing attack, which can be generalized across all robust watermarking schemes.
Figure 5: The relationship between $s/l$ and $z$. The data points are evaluated on Unigram using LLAMA-2-7B and 500500500 samples from OpenGen dataset.
Appendix D Proof of Theorem 2
Proof.
Recall that KGW and Unigram randomly split the tokens in the vocabulary into the green list and the red list. We consider greedy sampling, where the token with the highest (watermarked) probability is sampled. We have $n$ independent observations under different watermark keys. For each key, the token $\textbf{x}_{i}$ with the highest unwatermarked probability in the green list is $\gamma$. As long as $\textbf{x}_{i}$ is the green list, the greedy sampling will always yield $\textbf{x}_{i}$ since the watermarks add the same constant to all the tokens’ logits in the green list.
Thus, the probability that the most frequent token among these $n$ observations is $\textbf{x}_{i}$ is at least:
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}\gamma^{k}(1-\gamma)^{n-k},$
which is the probability that $\textbf{x}_{i}$ is in the green list for at least half of the $n$ keys.
For another token $\textbf{x}_{j}$ whose probability can exceed $\textbf{x}_{i}$, if $\textbf{x}_{j}$ is in the green list and $\textbf{x}_{i}$ is in the red list. Then if $\textbf{x}_{i}$ is in the green list for $k$ keys, the probability that $\textbf{x}_{j}$ is in the green list for at least $k$ keys among the other $n-k$ keys is:
$1-\sum_{m=0}^{k-1}\binom{n-k}{m}\gamma^{m}(1-\gamma)^{n-k-m}$
Consider we have $c$ such tokens having the potential to exceed $\textbf{x}_{i}$. Then at least one of the $c$ tokens is in the green list for at least $k$ keys among the other $n-k$ keys:
$1-\Bigl{(}\sum_{m=0}^{k-1}\binom{n-k}{m}\gamma^{m}(1-\gamma)^{n-k-m}\Bigr{)}^{c}$
Thus, with all the above analysis, we have that if there are $c$ tokens that have the potential to exceed the probability of the token with the highest unwatermarked probability (i.e., $\textbf{x}_{i}$), the probability that the most frequent token among the $n$ observations is the same as $\textbf{x}_{i}$ is:
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}\gamma^{k}(1-\gamma)^{n-k}\times\Biggl{(}1-\Bigl{(}\sum_{m=0}^{k-1}\binom{n-k}{m}\gamma^{m}(1-\gamma)^{n-k-m}\Bigr{)}^{c}\Biggr{)},$
which concludes the proof. ∎
Appendix E Probability Bound of Unwatermarked Token Estimation for Exp
In this section, we present and prove the probability bound of unwatermarked token estimation for the Exp watermark [KTHL23].
Theorem 3 (Probability bound of unwatermarked token estimation for Exp).
Suppose there are $n$ observations under different keys, the highest probability for the unwatermarked tokens is $p$. Then the probability that the most frequently appeared token among the $n$ observations is the same as the original unwatermarked token with the highest probability is:
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}p^{k}(1-p)^{n-k}$
Proof.
The proof of Theorem 3 is straightforward. As we have introduced in Appendix A, the Exp watermark employs the Gumbel trick sampling [Gum48] when embedding the watermark. Thus, the probability that we observe the token whose original unwatermarked probability is $p$ is exactly $p$ for each of the independent keys. Thus, if we make $n$ observations under different keys, then at least half of them yield the token with the highest original probability $p$:
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}p^{k}(1-p)^{n-k},$
which concludes the proof. ∎
Appendix F Algorithms of Attacks Exploiting the Detection API
In this section, we provide the detailed algorithm of the attacks exploiting the detection API as we have introduced in Sec. 6. Specifically, we present the algorithm for watermark-removal attack exploiting the detection API in Alg. 1 and the algorithm for spoofing attack exploiting the detection API in Alg. 2.
Algorithm 1 Watermark-removal attack exploiting the detection API.
  Input: Prompt $\textbf{x}_{\text{prompt}}$, watermarked LLM $M_{\text{wm}}$, detection API $f_{\text{detection}}$, maximum output token number $m\geq 2$
  Let $k\leftarrow 5$, $\textbf{x}_{1}\sim M_{\text{wm}}(\textbf{x}_{\text{prompt}})$
  for $t=2$ to $m$ do
  $(\textbf{x}_{t}^{1},\textbf{x}_{t}^{2},\cdots,\textbf{x}_{t}^{k}),(\textbf{p}_{t}^{1},\textbf{p}_{t}^{2},\cdots,\textbf{p}_{t}^{k})\leftarrow M_{\text{wm}}(\textbf{x}_{prompt}||\textbf{x}_{1}\cdots\textbf{x}_{t-1})$The watermarked LLM returns the top $k$ tokens and their corresponding probabilities in descending order.}
  for $i=1$ to $k$ do
   $d_{i}\leftarrow f_{\text{detection}}(\textbf{x}_{1}||\cdots||\textbf{x}_{t-1}||\textbf{x}_{t}^{i})$
  $d_{\text{min}}\leftarrow\min(d_{1},d_{2},\cdots,d_{k})$, $l_{\text{candidate}}\leftarrow\text{empty}$Get the detection score with the lowest confidence.}
  for $i=1$ to $k$ do
   if $d_{\text{min}}=d_{i}$ then
    $l_{\text{candidate}}\leftarrow l_{\text{candidate}}||\textbf{x}_{t}^{i}$Get all the tokens with the lowest detection confidence.}
  if $\textbf{x}_{t}^{1}\in l_{\text{candidate}}$ then
   $j\leftarrow 0$If the token with the highest probability (the first token) is in the list, output that token.}
  else
   $c\leftarrow 1$
   for $\textbf{x}_{t}^{i}\in l_{\text{candidate}}$ do
    $\textbf{p}_{t}^{i}\leftarrow\textbf{p}_{t}^{1}/c$Update the tokens’ probabilities that have lowest detection confidence scores.}
    $c\leftarrow c+1$
   $\textbf{p}_{t}^{1}\leftarrow 0$
   $j\leftarrow\text{Sample}(\textbf{p}_{t}^{1},\cdots,\textbf{p}_{t}^{k})$Sample the tokens according to the updated probabilities.}
  $\textbf{x}_{t}\leftarrow\textbf{x}_{t}^{j}$
  Return $\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{m}$
Algorithm 2 Spoofing attack exploiting the detection API.
  Input: Prompt $\textbf{x}_{\text{prompt}}$, local LLM $M$, detection API $f_{\text{detection}}$, maximum output token number $m$
  Let $k\leftarrow 3$
  for $t=1$ to $m$ do
  $(\textbf{x}_{t}^{1},\textbf{x}_{t}^{2},\cdots,\textbf{x}_{t}^{k}),(\textbf{p}_{t}^{1},\textbf{p}_{t}^{2},\cdots,\textbf{p}_{t}^{k})\leftarrow M(\textbf{x}_{prompt}||\textbf{x}_{1}\cdots\textbf{x}_{t-1})$The local LLM returns the top $k$ tokens and their corresponding probabilities in descending order.}
  for $i=1$ to $k$ do
   $d_{i}\leftarrow f_{\text{detection}}(\textbf{x}_{1}||\cdots||\textbf{x}_{t-1}||\textbf{x}_{t}^{i})$
  $j\leftarrow\arg\max(d_{1},d_{2},\cdots,d_{k})$Get the token resulting in the highest confidence.}
  $\textbf{x}_{t}\leftarrow\textbf{x}_{t}^{j}$
  Return $\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{m}$
Appendix G Additional Results of Watermark-Removal Attacks Exploiting Quality-Preserving Property
In this section, we provide more evaluation results of the watermark-removal attack exploiting the quality-preserving property (see Sec. 5) on all the three watermarks (KGW, Unigram, and Exp) and two models (Llama-2-7B and OPT-1.3B). The results are shown in Fig. 678910. For KGW watermark on OPT-1.3B model and Unigram watermark on Llama-2-7B and OPT-1.3B models, we have consistent observations with the KGW watermark on Llama-2-7B as we present in Sec. 5.1, demonstrating the effectiveness of our attacks. For the Exp watermark, our results in Fig. 7 and Fig. 10 also show the watermark can be easily removed using multiple queries to estimate the unwatermarked tokens.
For Exp watermark [KTHL23], the use of multiple watermark keys is inherent in its design, which is default to 256256256. Thus, Exp has the same unwatermarked detection results for various numbers of queries under different watermark keys. From the results in Fig. 7 and Fig. 10, we conclude that using $n=13$ queries under different keys, the resulting p-value is very close to that of the content without watermark and is significantly different from the watermarked p-value, which shows that we can effectively remove the watermark using 131313 queries for each token. We note that for Exp, the perplexity of watermarked content is significantly higher than that of unwatermarked content. This is primarily because Exp does not allow sampling in watermark embedding, which becomes a deterministic algorithm when the key is fixed. Conversely, our watermark-removal attack generates content with much lower perplexity, making it comparable to unwatermarked content when the query number under different keys exceeds 131313. This can be attributed to our attack functioning as a layer of random sampling. Unlike greedy sampling methods, we have a probability to sample the token with the highest unwatermarked probability (refer Sec. 4, Appendix D, and Appendix E).
The results of the three watermarks and two models prove that the watermark-removal attack exploiting the quality-preserving property using multiple keys can effectively eliminate the watermarks while maintaining high output quality. We anticipate that this attack can be generalized to all watermarking schemes that have quality-preserving properties.
(a) Z-Score.
(b) Perplexity.
Figure 6: Watermark-removal on Unigram watermark [ZALW24] and Llama-2-7B model with multiple watermark keys.
(a) P-Value.
(b) Perplexity.
Figure 7: Watermark-removal on Exp watermark [KTHL23] and Llama-2-7B model with multiple watermark keys.
(a) Z-Score.
(b) Perplexity.
Figure 8: Watermark-removal on KGW watermark [KGW+23a] and OPT-1.3B model with multiple watermark keys.
(a) Z-Score.
(b) Perplexity.
Figure 9: Watermark-removal on Unigram watermark [ZALW24] and OPT-1.3B model with multiple watermark keys.
(a) P-Value.
(b) Perplexity.
Figure 10: Watermark-removal on Exp watermark [KTHL23] and OPT-1.3B model with multiple watermark keys.
Appendix H Additional Results of Attacks Exploiting Detection APIs
[…]
(a) Z-Score/P-Value of removal.
(b) Perplexity of removal.
(c) Z-Score/P-Value of spoofing.
Figure 11: Attacks exploiting detection APIs on OPT-1.3B model.
wm-removal spoofing
ASR #queries ASR #queries
KGW 0.990.990.99 2.872.872.87 1.001.001.00 2.962.962.96
Unigram 0.770.770.77 3.253.253.25 1.001.001.00 2.972.972.97
Exp 0.860.860.86 2.072.072.07 0.930.930.93 2.922.922.92
Table 3: The attack success rate (ASR), and the average query numbers per token for the watermark-removal and spoofing attacks exploiting the detection API on OPT-1.3B model.
Appendix I Additional Results of DP Defense
We present additional evaluation results of our defense technique that enhances the watermark detection by utilizing the techniques of differential privacy (see Sec. 6). Consistent with Sec. 6.3, we evaluate the utility of the DP defense as well as its performance in mitigating the spoofing attack exploiting the detection API. The results are shown in Fig. 1213141516.
[…]
Our defense can be generalized to all LLM watermarking schemes. It allows us to substantially mitigate spoofing attacks exploiting the detection API while having a negligible impact on utility.
(a) Detection accuracy and spoofing attack success rate.
(b) Z-scores with/without DP and under multiple queries.
Figure 12: Evaluation of DP watermark detection on Unigram watermark and Llama-2-7B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) P-values with/without DP and under multiple queries.
Figure 13: Evaluation of DP watermark detection on Exp watermark and Llama-2-7B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) Z-scores with/without DP and under multiple queries.
Figure 14: Evaluation of DP watermark detection on KGW watermark and OPT-1.3B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) Z-scores with/without DP and under multiple queries.
Figure 15: Evaluation of DP watermark detection on Unigram watermark and OPT-1.3B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) P-values with/without DP and under multiple queries.
Figure 16: Evaluation of DP watermark detection on Exp watermark and OPT-1.3B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).